40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5351     Entry time: Wed Sep 7 00:01:23 2011
Author: Suresh 
Type: Update 
Category: IOO 
Subject: ITMY chamber ready for heavy doors 

[Jenne, Suresh]

We did the following things in the ITMY chamber today:

1) We tried to get the ITMY stuck again by adjusting the coil gains so that it goes into the orientation where it used to get stuck.  We (reassuringly) failed to get it stuck again.  This, as we came to know later, is because kiwamu had rotated the side OSEM such that the optic does not get stuck . However the OSEM beam is at about 30 deg to the vertical and the SD is sensitive to POS motion now resulting in the poorer separation of modes as noted by Jenne earlier (5439)

2) We checked the earthquake stops and repositioned two at the bottom (towards the AR side of the optic)  which we had backed out earlier.

3) We took pics of all the OSEMS.

4) Checked to see if there are any stray beams with an IR card.  There were none.

5) I obtained the max values of the OSEMS by misaligning the optic with the coil offsets.  These values are in good agreement with those on the wiki

OSEM     UL     UR     LR     LL      SD

Max      1.80    1.53   1.68   1.96    2.10

Current  0.97   0.79    0.83   0.97   1.02

 

We can close the heavy doors tomorrow morning.

ELOG V3.1.3-